OpenStudy (anonymous):

This is a Quadratic Equation, solve it by factoring please.Also please explain. 3x^2+22x+15=5(x+1)

OpenStudy (ybarrap):

3 x^2+22 x+15 = 5 (x+1) Expand out terms on the right hand side: 3 x^2+22 x+15 = 5 x+5 Subtract (5 x+5) from both sides: 3 x^2+17 x+10 = 0 Solve the quadratic equation by completing the square: Divide both sides by 3: x^2+(17 x)/3+10/3 = 0 Subtract 10/3 from both sides: x^2+(17 x)/3 = -10/3 Add 289/36 to both sides: x^2+(17 x)/3+289/36 = 169/36 Factor the left hand side: (x+17/6)^2 = 169/36 Take the square root of both sides: abs(x+17/6) = 13/6 Eliminate the absolute value: x+17/6 = -13/6 or x+17/6 = 13/6 Subtract 17/6 from both sides: x = -5 or x+17/6 = 13/6 Subtract 17/6 from both sides: x = -5 or x = -2/3

OpenStudy (anonymous):

that is as long as its 3x squared.

hero (hero):

3x^2 + 15x + 2x + 10 = 0 3x(x+ 5) + 2(x+5) = 0 (x+5)(3x+2) = 0 x = -5, -2/3

hero (hero):

Solved by factoring

OpenStudy (anonymous):

thnnks everyone